OpenStudy (anonymous):

log(10^6x^2) = a. 12 b. 6+2log(x) c. 12log(x) d. 6x^2

OpenStudy (anonymous):

log(1000000 x^2)

OpenStudy (anonymous):

That's not one of the choices though

OpenStudy (ranga):

log(A * B) = log(A) + log(B) log(A^n) = n * log(A) log(10^6x^2) = log(10^6) + log(x^2) = 6log(10) + 2log(x) = 6 + 2log(x)